Grade 12Physical Chemistry

A solution is prepared by adding 2 g of a substance A to 18 g of water. Calculate the mass per cent of the solute

sudhanshu
12 Years agoGrade 12

2 Answers

Gaurav
11 Years ago
Hello Student
Mass percentage of solute = mass of solute in g/total mass of solution
= 2/(18+2) = 2/20 = 0.1%
Sherlock
7 Years ago
Mass percentage = mass of solute / mass of         solution × 100 
= 2 / 2 + 18 × 100 = 10% you are welcome 
"They said answer should not be less than 100 characters so this doesn't belong to the solution"